OpenStudy (anonymous):

what are these two radicals simplified? 48 and 60

OpenStudy (jdoe0001):

\(\bf \sqrt{48}\sqrt{60}\quad ?\)

OpenStudy (anonymous):

yes that! ^^^

OpenStudy (jdoe0001):

let us simplify both numbers, notice all we'll be doing is factoring them, so-called "simplify" and you'd do the same for any really, same way  \(\bf \sqrt{48}\quad \sqrt{60}\\ \quad \\ \quad \\ \color{green}{48 = 2\times 2\times 2\times 2\times 3\implies 2^2\times 2^2\times 3\\ \quad \\ \implies (2^2)^2\times 3\\ \quad \\ 60 = 2\times 2\times 3\times 5\implies 2^2\times 3\times 5}\\ \quad \\ \sqrt{48}\quad \sqrt{60}\implies \sqrt{(2^2)^2\times 3}\quad \sqrt{2^2\times 3\times 5}\) can you take anything out of the radicals?